Evaluate the double integral sin(y^3)dA where R is the region enclosed by y=√x , y=2,y=0 .


1
Expert's answer
2021-02-25T05:06:51-0500

Given region is -

"f(x,y)=sin^3y" and region is "y=\sqrt{x},y=0,y=2"


So The integral is given by


The limits for double integral for "x= 0 \text{ to }x=y^2" and for "y=0 \text{ to } y=2"


="\int_0^2\int_0^{y^2}sin(y^3)dxdy"


="\int_0^2sin(y^3)[x]_0^{y^2}dy"


="\int_0^2sin(y^3)y^2dy"


let "y^3=t\implies 3y^2dy=dt\implies y^2dy=\dfrac{dt}{3}"


At "y=0, t=0 \text { and at } y=2,t=y^3=(2)^3=8"


"=\int_0^{8}\dfrac{sint}{3}dt"


"=\dfrac{1}{3}[-cost]_0^{8}"


"=\dfrac{1}{3}(cos0-cos64)"


"=\dfrac{1}{3}(1-cos8)"


"=\dfrac{1}{3}(1-0.964)"


"\\=\dfrac{0.036}{3}\\=0.012 \text{ sq. units }"